OpenStudy (anonymous):

WILL MEDAL OR FAN PLEASE ANSWER!! If Oscar the ostrich travels 84 miles in 2 hours on land, how fast does he travel? A. 43 mph B. 42 mph C. 40 mph D. 163 mph

jhonyy9 (jhonyy9):

84 miles in 2 hours x mile --- 1 hour ------------------ x = 84*1/2 = ? mile/hour

jhonyy9 (jhonyy9):

x=(84*1)/2 = ?

OpenStudy (anonymous):

ok so we dont know how many miles there is so thats x right?? and 84x1 is 84 divide by 2? equeals 42

OpenStudy (anonymous):

right?

jhonyy9 (jhonyy9):

yes - sure right

OpenStudy (anonymous):

ok thanks

OpenStudy (vhtran):

B is the right choice. mph stands for Miles Per Hour, since it gives you 2 hours just divide 84 by 2

OpenStudy (anonymous):

ok thank you im gonna close this now